Solve using the sustitution method:
11x-6y=494
x+7y=-23

Answer by longjonsilver(2297)

ok then. First --> look at my Lessons on Simultaneous equations. They should help you.

equn2, re-written, is x=-23-7y. We can now put this into eqn1, so...

11(-23-7y)-6y=494. Hopefully, you can follow this? This is the "substitution" mentioned in the question. Now just find y!...

-253-77y-6y=494
-253-83y=494
-83y = 747
y = 747/(-83) --> y=-9.

OK, now we have y, we need to put the fact that y is 9 into either eqn1 or eqn2 --> your choice (i would suggest you work with the easier looking one). You try. x should be about 40-ish :-).

jon